9) what is the value of ΔΗ for a reaction if K at 50°C is 40% of K at 37°C?
Answer #1

There is a relation between enthalpy temperature and rate constant known as Van't Hoff's equation

K1

Here

T1 = 50 0C = 323K

T2 = 37 0C = 310K

K1 =0.4K2

DeltaH = ?

R = gas constant =8.314 J / mol K

Putting the values in equation

K2 0.4K28.314 310 323

or

0.916 =ーー(1.298x10-4) 8.314

Therefore

AHー-58.67 k J/mol

ln(K2/0.4K2) = (ΔH / 8.314) (1/